The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by algebrapro18(249)

if x is equal to the length of the base of the triangle then 4x-8 is equal to the length of the two remaining sides. Now just solve the equation for x.

2(4x-8) + x = 29
8x - 16 + x = 29
9x - 16 = 29
9x = 45
x = 5

so the length of one of the legs is:
4(5)-8
20-8
12